# Remarks on Dirac–Bergmann algorithm, Dirac’s conjecture and the extended Hamiltonian

### v2

@@ 1 Introduction
Most physicists naturally discuss the Hamiltonian formalism in the context of quantization, and various alternative approaches exist for that purpose [11, 12, 13, 14]. Our focus here, however, is on Hamiltonian analysis merely as a tool for studying classical theories, for which, in our opinion, the Dirac–Bergmann recipe provides the most convenient framework. Although it lacks some of the elegance inherent in the Lagrangian formulation, this method is very useful in that it offers a clear iterative procedure: it consists of a sequence of logical steps that reliably leads one to the result [^1] We recommend this paper[15], which illustrates all the major steps of the algorithm through the example of a single finite-dimensional system.. The key advantage is that it allows one to rewrite complicated nonlinear systems of second-order equations as first-order equations for canonical pairs.
@@
This work builds on papers [16, 17, 18] that discuss subtle and controversial features of the Dirac–Bergmann approach. These features concern the relation between gauge generators and first-class constraints, as well as the extended Hamiltonian, which were introduced in Dirac’s Lectures on Quantum Mechanics[6]. We insist that one has to exercise care when using this recipe, as these subtleties can actually make the analysis of a system more obscure. The addition of secondary first-class constraints to the Hamiltonian modifies the content of the theory in a peculiar way that keeps the number of dynamical modes intact and makes all first-class constraints independent gauge generators. The latter is not true in the case of the total Hamiltonian, where the correct generators are certain combinations of primary and secondary first-class constraints [19, 18]. The preservation of the number of dynamical modes is usually mentioned as a justification of Dirac’s extended Hamiltonian. However, “constrained” does not mean “unphysical” [16, 17]. Even though the number of dynamical degrees of freedom remains the same, the system is not equivalent to the one we started with in terms of the canonical variables. Solutions that describe static configurations become parameterized by arbitrary Lagrange multipliers associated with the secondary first-class constraints in the extended Hamiltonian. Either we can accept this and obtain a strange modification of the original theory, or we can redefine the system by declaring that gauge invariant quantities are those that commute with all first-class constraints. [17, 18]. The first option means a loss of static solutions; for example, Coulomb’s law in electrodynamics becomes unphysical. The second is an effective enlargement of the system’s gauge freedom via a reparameterization of variables. It is always attainable, which saves Dirac’s proposals.
@@
In this paper, we show that the above statements hold for every system with secondary first-class constraints that satisfies reasonable assumptions, including Yang–Mills theory and General Relativity. Additionally, it is natural to re-examine some aspects of this formalism given the recent realization that the Dirac–Bergmann algorithm has issues when applied to certain models [17, 20].
@@ 2 A brief reminder of Dirac–Bergmann recipe
Let us say we have a Lagrangian with $n$ generalized coordinates $\{q_{i}\}$. In the standard procedure for transitioning to the Hamiltonian formalism, one introduces the generalized momenta $p^{i}=\frac{\partial L}{\partial\dot{q}_{i}}(q,\dot{q})$ and then defines the Hamiltonian via the Legendre transform with respect to the velocities, $H=\dot{q}_{i}p^{i}-L$ [^2] Repeated indices imply summation.. Variation of it together with the requirement of Lagrange equations yields Hamilton’s canonical equations. This recipe works for usual models considered in mechanics courses because the Hessian w.r.t. velocities, $\frac{\partial^{2}L}{\partial\dot{q}_{i},\partial\dot{q}_{j}}$}, is non-degenerate. If the Hessian is degenerate, however, the definitions of the momenta are no longer invertible in terms of velocities, and the above procedure must be modified. The Hessian maps variations of the velocities to variations of the momenta (assuming the coordinates are fixed parameters). If the rank of the Hessian is equal to nKn-K, then only nKn-K independent variations of velocities are in one-to-one correspondence with nKn-K independent variations of momenta [^3] We assume that the mapping $\dot{q}\mapsto p$, defined by definitions of momenta, is regular in some region of $(q,\dot{q})$ values, or in other words, the rank of the Hessian in this region remains constant. This condition ensures that the number of constraints remains constant when transitioning to the canonical description later on.. The same therefore holds for the velocities and momenta themselves [^4] We will refer to these variables as resolved, and to the remaining ones as unresolved.. The remaining $K$ velocity variations are arbitrary free variables, and the corresponding momenta are constrained to be functions of other canonical variables, since their variations are arbitrary linear combinations of the nKn-K resolved momenta variations. The $K$ constrained momenta impose restrictions on the allowed values of variables. These are called primary constraints. Their general form[19, 21] can be written as follows
@@
where $\{p^{A}\}$ denotes the nKn-K momenta that can be uniquely expressed through velocities $p^{A}=f^{A}(q,\dot{q}^{B})$. This means our system does not have access to the entire phase space $(q,p)$. Rather, it is constrained to a hypersurface defined by (1). Since the functions $f^{a}$ do not depend on the unresolved velocities while being the right-hand side of the definitions of momenta we can write $L(q,p^{A},\dot{q}_{b})=\dot{q}_{a}f^{a}(q,p^{A})+\tilde{L}(q,p^{A})$.
Making the Legendre transformation of the Lagrangian w.r.t. velocities we get
$H(q,p,\dot{q}_{b})=f^{-1}_{A}(q,p^{B})p^{A}-\tilde{L}(q,p^{B})+\dot{q}_{a}[p^{a}-f^{a}(q,p^{B})].$ (2)
The unresolved velocities remain arbitrary and can therefore be identified with Lagrange multipliers, $\dot{q}_{a}=\lambda_{a}$. The resulting Hamiltonian is called the total Hamiltonian. It can be rewritten as
@@
where $H_{can}(q,p^{B})=f^{-1}_{A}(q,p^{B})p^{A}-\tilde{L}(q,p^{B})$ is the canonical Hamiltonian [14]. It coincides with the Legendre transform of the Lagrangian w.r.t. the velocities, with the primary constraints (1) imposed.
@@
Classification of constraints. Constraints in a Hamiltonian system are not limited to just primary ones. To ensure that the primary constraints $\phi^{a}(q(t),p(t))=0$ hold at all times, their time derivatives must vanish: $\dot{\phi}^{a}=\{\phi^{a},H_{T}\}=0$. These are the consistency conditions. They may either be satisfied identically, lead to restrictions on the Lagrange multipliers, or give rise to new constraints $\chi^{u}(q,p)=0$ [^5] In his lectures [6], Dirac introduces the concepts of weak ($\approx$) and strong (==) equality. A weak equality holds on the hypersurface defined by all constraints, while a strong equality holds on the entire phase space. In this work, we will not introduce a separate notation for equality on the constraint surface., which are called secondary constraints. Preservation of secondary constraints in time may in turn produce further constraints. These are also called secondary [^6] Constraints are usually viewed as restrictions on initial data, since they remain satisfied over time if imposed initially. In this work, however, we treat primary constraints as additional equations within the Hamiltonian system (4). In this approach, secondary constraints follow as consequences of these equations.. Secondary constraints further reduce the hypersurface on which the system can reside to a hypersurface of smaller dimension.
@@
Gauge symmetry in constrained Hamiltonian systems. All gauge systems are constrained systems due to the fact that gauge freedom necessarily leads to degeneracy of the Lagrangian. Gauge systems also stand out among all Hamiltonian systems with constraints, since they inevitably possess first-class constraints. The intuition is as follows: requiring a first-class constraint to hold at all times cannot fix any of the Lagrange multipliers in the total Hamiltonian, because the Poisson bracket of this constraint with the other constraints in the Hamiltonian vanishes. As a result, some of the Lagrange multipliers remain completely arbitrary, and this arbitrariness in the equations of motion is the gauge freedom: replacing one set of Lagrange multipliers with another in equations (4) amounts to a gauge transformation. However, the question of how first-class constraints are related to the generators of such transformations is rather subtle.
The standard answer in the literature is the so-called Dirac’s conjecture [6, 14]. It states that all first-class constraints, both primary and secondary, are independent generators of gauge transformations. In other words, for any quantity $g(q,p)$ its variation under a gauge transformation has the form
@@
Following this conjecture, Dirac suggested adding all first-class constraints to the Hamiltonian. The resulting Hamiltonian is called the extended Hamiltonian $H_{E}=H_{can}+\lambda_{\rho}\Omega^{\rho}$. In Dirac’s logic, adding secondary constraints to the Hamiltonian cannot change the physical content, since they generate transformations that do not change the physical state.
Remark. All statements in this section are trivially generalized to field theories, with the only caveat being that spatial coordinates play the role of an index labeling the generalized coordinates. In other words, field theories possess an infinite number of degrees of freedom.
@@ 3 A simple toy model
  • We assume that the system’s equations of motion are of order not higher than second in time. The total number of modes $N_{tot}$ denotes the total number of real-valued variables in Lagrangian formulation of the system.
@@
  • We define the number of dynamical modes $N_{dyn}$ as half of the minimal number of Cauchy data values required to make the evolution of the non-pure gauge variables unique. This definition is motivated by the analogy with classical mechanics: for a particle one must specify two initial data, its position and its velocity. The requirement of minimality of $N_{dyn}$ is necessary, since often there is no unambiguous way to decide which variables should be regarded as pure gauge and which should not according to the definition above. We resolve this by identifying as pure gauge those variables that minimize $N_{dyn}$.
  • We define the number of constrained modes as $N_{con}=N_{tot}-N_{dyn}-N_{gauge}$. Constrained modes correspond to variables that are uniquely determined by the equations of motion, so no initial conditions needed to be specified for them. We call them physical, since they follow directly from the equations of motion once the pure gauge variables are taken into account. For the field theories we are interested in, they correspond to static configurations.
Let us stress one more point. In the common view, constrained modes are often regarded as unphysical. This is usually motivated by the fact that one can always, at least in principle, solve all constraints and express the corresponding variables in terms of the remaining ones. This is true. However, in electrodynamics and gravity, the constrained modes encode static field-configurations – Coulomb’s and Newton’s laws. Calling them unphysical is therefore a rather odd claim. Regardless of terminology, this physical information is still present in the equations, even after the constraints are solved, and it must be taken into account when analyzing the system.
@@
Situations of this type are often referred to as “gauge hits twice” or “gauge strikes twice” [14, 16] in the jargon. The peculiarity is that the presence of gauge symmetry leads to two first-class Hamiltonian constraints for every independent symmetry (a primary one and the corresponding secondary one), so that two dynamical modes are eliminated at once: one becomes pure gauge, and the other becomes constrained because of the pair of constraints. In our case the counting gives $N_{dyn}=0$, $N_{gauge}=1$, $N_{con}=1$.
@@ 3.1 Castellani Generators
The proof of the theorem is based on applying perturbation theory in the gauge parameters to the equations of motion (4) [19, 18]. Unlike Dirac’s argument, time does not enter the proof as an expansion parameter. To obtain the chains (16), one considers the variation of equations (4) to first order in the small parameters $\delta q$ and $\delta p$ and requires that all terms distinguishing the transformed equations from the original ones vanish.
The chain (16) implies that all $G_{k}$ are first-class constraints. Moreover, each $G_{k>0}$ includes secondary first-class constraints ($G_{k}$ is not necessarily a secondary constraint itself, since $G_{k}$ may also contain a linear combination of primary first-class constraints) [^10] Nevertheless, in general the generators do not include all first-class constraints of the theory. An exception arises for constraints of type $\chi^{n}$ and for those that follow from them via the consistency conditions. By constraints of type $\chi^{n}$ we mean constraints with the structure $\chi^{n>1}$, where $\chi=0$ is a constraint that vanishes to first order on the constraint surface..
@@
Let us note once again that the Castellani generator $G_{C}$ is a gauge symmetry generator in the standard sense – it generates a local transformation that maps any solution of the equations of motion to another solution describing the same physical state. However, if we restrict ourselves to transformations that preserve the physical state at a fixed instant of time, then Dirac’s conjecture does work, since at that instant all time derivatives of the parameter $\varepsilon$ become independent. [^11] One may notice that Dirac’s conjecture is satisfied in the usual notion of gauge symmetry in the trivial case when the system has only primary first-class constraints. [^12] An exception, again, is provided by constraints of type $\chi^{n}$. A simple example is the system $L=\frac{1}{2}e^{y}\dot{x}^{2}$ discussed in [14]. The system has two first-class constraints, yet only the primary one generates transformations of the spurious variable $y$. At the same time, it is hard to interpret these transformations as a genuine gauge symmetry, since they change the action [16]. Geometrically, such transformations can be viewed as symmetries of the phase space, foliated into layers – equivalence classes. Two points belong to the same class if they are related by a transformation generated by the Dirac generator $G_{D}$. This distinction, which is crucial in our view, is not always made explicit in the literature. As a result, Dirac’s conjecture is often formulated using different underlying notions of gauge symmetry. For example, [24] adopts the standard interpretation, whereas [14] considers gauge transformations only at a fixed instant of time. In this paper, we examine the validity of the conjecture in both formulations.

### v3

@@ 1 Introduction
Most physicists naturally discuss the Hamiltonian formalism in the context of quantization, and various alternative approaches exist for that purpose [11, 12, 13, 14]. Our focus here, however, is on Hamiltonian analysis merely as a tool for studying classical theories, for which, in our opinion, the Dirac–Bergmann algorithm provides the most convenient framework. Although it lacks some of the elegance inherent in the Lagrangian formulation, this method is very useful in that it offers a clear iterative procedure: it consists of a sequence of logical steps that reliably leads one to the result [^1] We recommend this paper[15], which illustrates all the major steps of the algorithm through the example of a single finite-dimensional system.. The key advantage is that it allows one to rewrite complicated nonlinear systems of second-order equations as first-order equations for canonical pairs.
@@
This work builds on papers [16, 17, 18] that discuss subtle and controversial features of the Dirac–Bergmann approach. These features concern the relation between gauge generators and first-class constraints, as well as the extended Hamiltonian, which were introduced in Dirac’s Lectures on Quantum Mechanics[6]. We insist that one must exercise care when using this recipe, as these subtleties can actually make the analysis of a system more obscure. The addition of secondary first-class constraints to the Hamiltonian modifies the content of the theory in a peculiar way that keeps the number of dynamical modes intact and makes all first-class constraints independent gauge generators. The latter is not true in the case of the total Hamiltonian, where the correct generators are certain combinations of primary and secondary first-class constraints [19, 18]. The preservation of the number of dynamical modes is usually mentioned as a justification of Dirac’s extended Hamiltonian. However, constrained modes are not unphysical [16, 17]. Even though the number of dynamical degrees of freedom remains the same, the system is not equivalent to the one we started with in terms of the original canonical variables. Solutions that describe static configurations become parameterized by arbitrary Lagrange multipliers associated with the secondary first-class constraints in the extended Hamiltonian. Either we can accept this and obtain a strange modification of the original theory, or we can redefine the system by declaring that gauge invariant quantities are those that commute with all first-class constraints [17, 18]. The first option means a loss of static solutions; for example, Coulomb’s law in electrodynamics becomes unphysical. The second is an effective enlargement of the system’s gauge freedom via a reparameterization of variables. It is always attainable, which saves Dirac’s proposals.
@@
In this paper, we show that the above statements hold for every system with secondary first-class constraints that satisfies reasonable assumptions, including Yang–Mills theory and general relativity. Additionally, it is natural to re-examine some aspects of the Dirac-Bergmann recipe, given the recent realization that it has issues when applied to certain models [17, 20].
@@ 2 A brief reminder of Dirac–Bergmann recipe
Let us say we have a Lagrangian with $n$ generalized coordinates $\{q_{i}\}$. In the standard procedure for transitioning to the Hamiltonian formalism, one introduces the generalized momenta $p^{i}=\frac{\partial L}{\partial\dot{q}_{i}}(q,\dot{q})$ and then defines the Hamiltonian via the Legendre transform with respect to the velocities, $H=\dot{q}_{i}p^{i}-L$ [^2] Repeated indices imply summation.. Varying the Hamiltonian, together with the requirement of the Lagrange equations, yields Hamilton’s canonical equations. This recipe works for the usual models considered in classical mechanics courses because the Hessian of the Lagrangian with respect to velocities, $\frac{\partial^{2}L}{\partial\dot{q}_{i},\partial\dot{q}_{j}}$}, is non-degenerate. If degeneracy is present, however, the definitions of the momenta are no longer invertible in terms of velocities, and the above procedure must be modified. The Hessian maps variations of the velocities to variations of the momenta (assuming the coordinates are fixed parameters). If the rank of the Hessian is equal to nKn-K, then only nKn-K independent variations of velocities are in one-to-one correspondence with nKn-K independent variations of momenta [^3] We assume that the mapping $\dot{q}\mapsto p$, given by the definitions of momenta, is regular in some region of $(q,\dot{q})$ values, or in other words, the rank of the Hessian in this region remains constant. This condition ensures that the number of constraints remains constant when transitioning to the canonical description later on. The Dirac-Bergmann approach is based on this crucial assumption.. Consequently, the same holds for the corresponding velocities and momenta themselves. We will refer to these as resolved variables, and to the remaining ones as unresolved.
The unresolved $K$ velocities are arbitrary free parameters, and their associated momenta are constrained to be functions of other canonical variables. The $K$ constrained momenta impose restrictions on the accessible states of the system. These are called primary constraints. Their general form[19, 21] can be written as follows
@@
where $\{p^{A}\}$ denotes the nKn-K momenta that can be uniquely expressed through the resolved velocities $p^{A}=f^{A}(q,\dot{q}_{B},\dot{q}_{a})$. This means our system does not have access to the entire phase space $(q,p)$. Rather, it is constrained to a hypersurface defined by (1).
Performing the Legendre transformation of the Lagrangian $L(q,\dot{q}_{A},\dot{q}_{b})$ with respect to velocities, we obtain
$H(q,p,\dot{q}_{b})=f^{-1}_{A}(q,p^{B},\dot{q}_{b})p^{A}-L(q,f^{-1}_{B}(q,p^{C},\dot{q}_{a}),\dot{q}_{b})+\dot{q}_{a}f^{a}(q,p^{B})+\dot{q}_{a}[p^{a}-f^{a}(q,p^{B})],$ (2)
where $f^{-1}_{A}(q,p^{B},\dot{q}_{a})$ is the inverse mapping from $p^{A}$ to $\dot{q}_{A}$, with the variables $(q,\dot{q}_{a})$ treated as fixed parameters. The unresolved velocities remain arbitrary and can therefore be identified with Lagrange multipliers, $\dot{q}_{a}=\lambda_{a}$. The resulting Hamiltonian is called the total Hamiltonian. It can be rewritten as
@@
where $H_{can}(q,p^{B})=f^{-1}_{A}(q,p^{B},\dot{q}_{b})p^{A}-L(q,f^{-1}_{B}(q,p^{C},\dot{q}_{a}),\dot{q}_{b})+\dot{q}_{a}f^{a}(q,p^{B})$ is the canonical Hamiltonian [14]. It coincides with the Legendre transform of the Lagrangian with respect to the velocities, with the primary constraints (1) imposed. One can straightforwardly verify that $H_{can}$ does not depend on the unresolved velocities.
@@
Classification of constraints. Constraints in a Hamiltonian system are not limited to just primary ones. To ensure that the primary constraints $\phi^{a}(q(t),p(t))=0$ hold at all times, their time derivatives must vanish: $\dot{\phi}^{a}=\{\phi^{a},H_{T}\}=0$. These are the consistency conditions [14]. They may either be satisfied identically, lead to restrictions on the Lagrange multipliers in the form of a system of linear equations, or give rise to new constraints $\chi^{u}(q,p)=0$ [^4] In his lectures [6], Dirac introduces the concepts of weak ($\approx$) and strong (==) equality. A weak equality holds on the hypersurface defined by all constraints, while a strong equality holds on the entire phase space. In this work, we will not introduce a separate notation for equality on the constraint surface., which are called secondary constraints. The consistency conditions for the secondary constraints may, in turn, give rise to tertiary or higher-order ones. All such subsequent constraints are also referred to as secondary [^5] Constraints are usually viewed as restrictions on initial data, since they remain satisfied over time if imposed initially. In this work, however, we treat primary constraints as additional equations within the system (4). In this approach, secondary constraints follow as consequences of these equations.. Secondary constraints further reduce the hypersurface on which the system can reside to a hypersurface of smaller dimension.
@@
Gauge symmetry in constrained Hamiltonian systems. Gauge systems stand out among all Hamiltonian constrained systems since they inevitably possess first-class constraints. The intuition is as follows: requiring a first-class constraint to hold at all times cannot fix any of the Lagrange multipliers in the total Hamiltonian because the Poisson bracket of this constraint with the other constraints in the Hamiltonian vanishes. As a result, some of the Lagrange multipliers remain completely arbitrary, and this arbitrariness in the equations of motion is the gauge freedom: replacing one set of Lagrange multipliers with another in equations (4) amounts to a gauge transformation. However, the question of how first-class constraints are related to the generators of such transformations is rather subtle.
The standard answer in the literature is the so-called Dirac’s conjecture [6, 14]. It states that all first-class constraints, both primary and secondary, are independent generators of gauge transformations. In other words, for any quantity $g(q,p)$, its variation under a gauge transformation has the form
@@
Following this conjecture, Dirac suggested adding all first-class constraints to the Hamiltonian. The resulting Hamiltonian is called the extended Hamiltonian $H_{E}=H_{can}+\lambda_{\rho}\Omega^{\rho}$. In Dirac’s logic, adding secondary constraints to the Hamiltonian cannot change the physical content since they generate transformations that do not change the physical state.
Remark. All statements in this section generalize to field theories, with the caveat that spatial coordinates play the role of an index labeling the generalized coordinates. In other words, field theories can be regarded as mechanical systems with an infinite number of degrees of freedom. Consequently, the standard recipe remains valid for well-behaved theories. However, we must note that there are prominent cases when the Dirac-Bergmann algorithm is known to be problematic, such as $f(\mathbb{Q})$ gravity models [20], where the consistency conditions yield a system of first-order partial differential equations for the Lagrange multipliers instead of the usual system of linear algebraic equations; the former generally lacks unique solutions. This leads the Dirac-Bergmann algorithm to produce an incorrect number of modes [^6] An alternative method of mode counting [22] might be of more use in such instances. Additionally, one can encounter models that possess singular loci in their phase spaces, making the number of dynamical modes ill-defined [17]. How to apply the algorithm to such systems is often unclear. We shall put these exotic examples aside in this paper, however, as we are only interested in exploring questions concerning the gauge generators and the extended Hamiltonian in the standard theory of constrained Hamiltonian systems.
@@ 3 A simple toy model
  • We assume that the system’s equations of motion are of order not higher than second in time. The total number of modes $N_{tot}$ denotes the total number of real-valued variables in the Lagrangian formulation of the system.
@@
  • We define the number of dynamical modes $N_{dyn}$ as half of the minimal number of Cauchy data values required to make the evolution of the non-pure gauge variables unique. This definition is motivated by the analogy with classical mechanics: for a particle, one must specify two initial data values: its position and its velocity. The requirement of minimality of $N_{dyn}$ is necessary, since there is often no unambiguous way to decide which variables should be regarded as pure gauge and which should not according to the definition above. We resolve this by identifying as pure gauge those variables that minimize $N_{dyn}$.
  • We define the number of constrained modes as $N_{con}=N_{tot}-N_{dyn}-N_{gauge}$. Constrained modes correspond to variables that are uniquely determined by the equations of motion, so no initial conditions needed to be specified for them. We call them physical, since they follow directly from the equations of motion once the pure gauge variables are taken into account.
Let us stress one more point. In the common view, constrained modes are often regarded as unphysical. This is usually motivated by the fact that one can always, at least in principle, solve all constraints and express the corresponding variables in terms of the remaining ones. This is true. However, in electrodynamics and gravity, the constrained modes encode static field-configurations – Coulomb’s and Newton’s laws. Calling them unphysical is therefore a rather odd claim. Nevertheless, regardless of terminology, this physical information is still present in the equations, even after the constraints are solved, and it must be taken into account when analyzing the system.
@@
Situations of this type are often referred to as “gauge hits twice” or “gauge strikes twice” [14, 16] in the jargon. The peculiarity is that the presence of gauge symmetry leads to two first-class Hamiltonian constraints for every independent symmetry (a primary one and the corresponding secondary one), so that two dynamical modes are eliminated at once: one becomes pure gauge, and the other becomes constrained because of the pair of constraints. In our case, the counting gives $N_{dyn}=0$, $N_{gauge}=1$, $N_{con}=1$.
@@ 3.1 Castellani Generators
The proof of the theorem is based on applying perturbation theory in the gauge parameters to the equations of motion (4) [19, 18]. Unlike in Dirac’s argument, time does not enter the proof as an expansion parameter. To obtain the chains (16), one considers the variation of the equations (4) to first order in the small parameters $\delta q$ and $\delta p$ and requires that all terms distinguishing the transformed equations from the original ones vanish. A thorough analysis of the expressions for local symmetries in terms of constraints can also be found in [25].
The chain (16) implies that all $G_{k}$ are first-class constraints. Moreover, each $G_{k>0}$ includes secondary first-class constraints ($G_{k}$ is not necessarily a secondary constraint itself, since $G_{k}$ may also contain a linear combination of primary first-class constraints) [^10] Nevertheless, in general, the generators do not include all first-class constraints of the theory. An exception arises for constraints of type $\chi^{n}$ and for those that follow from them via the consistency conditions. By constraints of type $\chi^{n}$, we mean those with the structure $\chi^{n>1}$, where $\chi=0$ is a regular constraint, i.e., its differential is non-vanishing on the constraint surface..
@@
Let us note once again that the Castellani generator $G_{C}$ is a gauge symmetry generator in the standard sense – it generates a local transformation that maps any solution of the equations of motion to another solution describing the same physical state. However, if we restrict ourselves to transformations that preserve the physical state at a fixed instant of time, then Dirac’s conjecture does work, since at that instant all time derivatives of the parameter $\varepsilon$ become independent. [^11] One may notice that Dirac’s conjecture is satisfied in the usual notion of gauge symmetry in the trivial case when the system has only primary first-class constraints. [^12] An exception, again, is provided by constraints of type $\chi^{n}$. A simple example is the system $L=\frac{1}{2}e^{y}\dot{x}^{2}$ discussed in [14]. The system has two first-class constraints, yet only the primary one generates transformations of the spurious variable $y$. At the same time, it is hard to interpret these transformations as a genuine gauge symmetry, since they change the action [16]. Geometrically, such transformations can be viewed as symmetries of the phase space, foliated into layers – equivalence classes. Two points belong to the same class if they are related by a transformation generated by the Dirac generator $G_{D}$. This distinction, which is crucial in our view, is not always made explicit in the literature. As a result, Dirac’s conjecture is often formulated using different underlying notions of gauge symmetry. For example, [26, 25] adopts the standard interpretation, whereas [14] considers gauge transformations only at a fixed instant of time. In this paper, we examine the validity of the conjecture in both formulations.